# The Curse of Depth in Large Language Models

### v1

@@ 4 LayerNorm Scaling
$\mathbf{h}^{(l)}=\text{LayerNorm}(\mathbf{h}^{(\ell)})\times\frac{1}{\sqrt{\ell}},$ (12)
where $\ell\in\{1,2,\dots,L\}$. This scaling prevents excessive variance growth with depth, addressing a key limitation of Pre-LN. Unlike Mix-LN, which stabilizes gradients in deeper layers but suffers from training instability caused by Post-LN (Nguyen and Salazar, 2019; Wang et al., 2024), LayerNorm Scaling preserves the stability advantages of Pre-LN while enhancing the contribution of deeper layers to representation learning. Applying LayerNorm Scaling leads to a notable reduction of layerwise output variance, resulting in lower training loss and faster convergence than vanilla Pre-LN. Moreover, compared with previous LayerNorm variants (Li et al., 2024b; Liu et al., 2020), LayerNorm Scaling is hyperparameter-free, easy to implement, and does not introduce additional learnable parameters, making it computationally efficient and readily applicable to existing Transformer architectures.
@@ 4.1 Theoretical Analysis of LayerNorm Scaling
$\sigma^{2}_{x_{\ell+1}}=\sigma_{x_{\ell}}^{2}\Theta(1+\frac{1}{\sqrt{\ell}\sigma_{x_{\ell}}}),$ (13)
@@
where $\epsilon$ is a small number with $40<\epsilon\leq 1/$ 4.
From Lemma 2, we can conclude that our scaling method effectively slows the growth of the variance upper bound, reducing it from exponential to polynomial growth. Specifically, it limits the upper bound to a quadratic rate instead of an exponential one. Based on Theorem 1, after scaling, we obtain the following:
Theorem 2.
@@
By comparing Theorem 1 (before scaling) with Theorem 17 (after scaling), we observe a substantial reduction in the upper bound of variance. Specifically, it decreases from exponential growth $\Theta(\exp(L))$ to at most quadratic growth $\Theta(L^{2})$. In fact, this growth is even slower than quadratic expansion, as it follows $\Theta(L^{(2-\epsilon)})$ for some small $\epsilon>0$.
@@ 5.1 LLM Pre-training
To evaluate the effectiveness of LayerNorm Scaling, we follow the experimental setup of Li et al. (2024b), using the same model configurations and training conditions to compare it with widely used normalization techniques, including Post-LN (Nguyen and Salazar, 2019), DeepNorm (Wang et al., 2024), and Pre-LN (Dai et al., 2019). In line with Lialin et al. (2023) and Zhao et al. (2024), we conduct experiments using LLaMA-based architectures with model sizes of 130M, 250M, 350M, and 1B parameters, ensuring consistency in architecture and training settings.
Table 1: Perplexity (↓) comparison of various layer normalization methods across various LLaMA sizes.
@@
The architecture incorporates RMSNorm (Shazeer, 2020) and SwiGLU activations (Zhang and Sennrich, 2019), which are applied consistently across all model sizes and normalization methods. For optimization, we use the Adam optimizer (Kingma, 2015) and adopt size-specific learning rates: $1\times 10^{-3}$ for models up to 350M parameters, and $5\times 10^{-4}$ for the 1B parameter model. All models share the same architecture, hyperparameters, and training schedule, with the only difference being the choice of normalization method. Unlike Mix-LN (Li et al., 2024b), which introduces an additional hyperparameter $\alpha$ manually set to 0.25, LayerNorm Scaling requires no extra hyperparameters, making it simpler to implement. Table 1 shows that LayerNorm Scaling consistently outperforms other normalization methods across different model sizes. While DeepNorm performs comparably to Pre-LN on smaller models, it struggles with larger architectures like LLaMA-1B, showing signs of instability and divergence in loss values. Similarly, Mix-LN outperforms Pre-LN in smaller models but faces convergence issues with LLaMA-350M, indicating its sensitivity to architecture design and hyperparameter tuning due to the introduction of Post-LN. Notably, Mix-LN was originally evaluated on LLaMA-1B with 50,000 steps (Li et al., 2024b), while our setting extends training to 100,000 steps, where Mix-LN fails to converge, highlighting its instability in large-scale settings caused by the usage of Post-LN.
In contrast, LayerNorm Scaling solves the Curse of Depth without compromising the training stability thanks to its simplicity. LayerNorm Scaling achieves the lowest perplexity across all tested model sizes, showing stable performance improvements over existing methods. For instance, on LLaMA-130M and LLaMA-1B, LayerNorm Scaling reduces perplexity by 0.97 and 1.31, respectively, compared to Pre-LN. Notably, LayerNorm Scaling maintains stable training dynamics for LLaMA-1B, a model size where Mix-LN fails to converge. These findings demonstrate that LayerNorm Scaling provides a robust and computationally efficient normalization strategy, enhancing large-scale language model training without additional implementation complexity.
Comparison with Other Layer Normalization. In addition, we conducted comparisons using LLaMA-130M to evaluate LayerNorm Scaling against recently proposed normalization methods, including Admin (Liu et al., 2020), Sandwich-LN (Ding et al., 2021), and Group-LN (Wu and He, 2018; Ma et al., 2024). Table 2 shows that Admin and Group-LN degrade performance. Sandwich-LN slightly outperforms Pre-LN. Both Mix-LN and LayerNorm Scaling improve over Pre-LN by good margins. However, Mix-LN fails to reduce perplexity under 26, falling short of LayerNorm Scaling.
Table 2: Comparison against other normalization methods on LLaMA-130M. Perplexity (↓) is reported.
Pre-LN Admin Group-LN Sandwich-LN Mix-LN LayerNorm Scaling
26.73 27.91 28.01 26.51 26.07 25.76
@@ 5.2 Supervised Fine-tuning
We believe that LayerNorm Scaling allows deeper layers in LLMs to contribute more effectively during supervised fine-tuning by alleviating gradient vanishing associated with increasing depth. Compared to models trained with Pre-LN, the deeper layers with LayerNorm Scaling maintain stable output variance, preventing uncontrolled growth and ensuring effective feature representation. As a result, deeper layers contribute more effectively to feature transformation, enhancing representation learning and improving generalization on complex downstream tasks.
Table 3: Fine-tuning performance ($\uparrow$) of LLaMA with various normalizations.
@@
To verify this, we follow the fine-tuning methodologies in Li et al. (2024b) and Li et al. (2024a), applying the same optimization settings as pre-training. We fine-tune models from Section 5.1 on the Commonsense170K dataset (Hu et al., (Hu et al., 2023)) across eight downstream tasks. The results, presented in Table 3, demonstrate that LayerNorm Scaling consistently surpasses other normalization techniques in all evaluated datasets. For the LLaMA-250M model, LayerNorm Scaling improves average performance by1.80% and achieves a 3.56% gain on ARC-e compared to Mix-LN. Similar trends are observed with the LLaMA-1B model, where LayerNorm Scaling outperforms Pre-LN, Post-LN, Mix-LN, and DeepNorm on seven out of eight tasks, with an average gain of 1.86% over the best baseline. These results confirm that LayerNorm Scaling, by improving gradient flow and deep-layer representation quality, achieves better fine-tuning performance, demonstrating robustness and enhanced generalization on diverse downstream tasks.
5.3 LayerNorm Scaling Reduces Output Variance
As LayerNorm Scaling aims to reduce output variance, we validate this by comparing it with two scaling approaches: LayerScale (Touvron et al., 2021) and Scaled Initialization (Shoeybi et al., 2020). LayerScale applies per-channel weighting using a diagonal matrix, diag$(\lambda_{1},\ldots,\lambda_{d})$, where each weight $\lambda_{i}$ is initialized to a small value (e.g., $\lambda_{i}=\epsilon$). Unlike LayerNorm Scaling, LayerScale learns the scaling factors automatically, which does not necessarily induce a down-scaling effect. Scaled Initialization scales the initialization of $W_{0}$ and $W_{2}$ to small values by $\frac{1}{\sqrt{2L}}$ where $L$ is the total number of transformer layers. Since scaling is applied only at initialization, we argue that Scaled Initialization may not effectively reduce variance throughout training. We further verify this in Figure 1, where we can see the output variance of Scaled Initialization is as large as Pre-LN. Table 4 presents the results of LLaMA-130M and LLaMA-250M. First, we can see that LayerScale degrades performance. While Scaled Initialization slightly improves over Pre-LN, it falls short of LayerNorm Scaling and the gap becomes larger for the larger model.
@@ 5.3 LayerNorm Scaling Reduces Output Variance
Table 4: Comparison against other scaling methods.
Perplexity ($\downarrow$) LLaMA-130M LLaMA-250M
@@
+ LayerScale 27.93 23.45
+ Scaled Initialization 26.04 20.98
@@
5.4 Enhancing Deep Layers with LayerNorm Scaling
Figure 4: Performance drop of layer pruning on LLaMA-130M. LayerNorm Scaling enables deep layers to make a meaningful contribution to the model.
To evaluate how LayerNorm Scaling improves deep layer effectiveness, we conduct a layer pruning experiment on LLaMA-130M, systematically removing individual layers and measuring the performance drop ($\Delta P^{(\ell)}$) on the ARC-e benchmark (Clark et al., 2018). Figure 4 compares the pruning effects between standard Pre-LN and LayerNorm Scaling. In the Pre-LN, removing deep layers results in minimal performance degradation, indicating their limited role in representation learning. In contrast, with LayerNorm Scaling, pruning deeper layers leads to a more pronounced drop, suggesting they now play a more active role in learning. While early layers remain critical in both models, the performance degradation in the LayerNorm Scaling is more evenly distributed across layers, reflecting a more balanced learning process. These findings confirm that LayerNorm Scaling mitigates the Curse of Depth by ensuring deeper layers contribute effectively to training.

### v2

@@ 4 LayerNorm Scaling (LNS)
$\mathbf{h}^{(\ell)}=\text{LayerNorm}(\mathbf{h}^{(\ell)})\times\frac{1}{\sqrt{\ell}},$ (12)
where $\ell\in\{1,2,\dots,L\}$. This scaling prevents excessive variance growth with depth, addressing a key limitation of Pre-LN. Unlike Mix-LN, which stabilizes gradients in deeper layers but suffers from training instability caused by Post-LN (Nguyen and Salazar, 2019; Wang et al., 2024), LayerNorm Scaling preserves the stability advantages of Pre-LN while enhancing the contribution of deeper layers to representation learning. Applying LayerNorm Scaling leads to a notable reduction of layerwise output variance as shown in Figure 4, resulting in a lower training loss. Moreover, compared with previous LayerNorm variants (Li et al., 2024b; Liu et al., 2020), LayerNorm Scaling is hyperparameter-free, easy to implement, and does not introduce additional learnable parameters, making it computationally efficient and readily applicable to existing Transformer architectures.
@@ 4.1 Theoretical Analysis of LayerNorm Scaling
$\sigma^{2}_{x_{\ell}}=\sigma_{x_{1}}^{2}\Theta\Big{(}\prod^{\ell-1}_{k=1}\Big{(}1+\frac{1}{\sqrt{k}\sigma_{x_{k}}}\Big{)}\Big{)},$ (13)
@@
where $\epsilon$ is a small number with $11/2\leq\epsilon<$ 1.
From Lemma 4.1, we can conclude that our scaling method effectively slows the growth of the variance upper bound, reducing it from exponential to polynomial growth. Specifically, it limits the upper bound to a quadratic rate instead of an exponential one. Based on Theorem 3.3, after scaling, we obtain the following:
Theorem 4.2.
@@
By comparing Theorem 3.3 (before scaling) with Theorem 17 (after scaling), we observe a substantial reduction in the upper bound of variance. Specifically, it decreases from exponential growth $\Theta(\exp(L))$ to at most quadratic growth $\Theta(L^{2})$. In fact, this growth is even slower than quadratic expansion, as it follows $\Theta(L^{(2-\epsilon)})$ for some small $\epsilon>0$.
@@
In addition to making the deeper layers more effective, our variance-scaling approach can also reduce sudden spikes in the loss landscape during training. Based on Takase et al. (2023b)’s work, We formalize this in the following theorem Theorem 4.3, which gives a rigorous upper bound on the gradient norm with respect to the attention parameters.
Theorem 4.3.
For the Pre-Ln transformers with weight $W_{1}$ on its first layer’s query projection. Then the $L$-layer backpropagated gradient norm with respect to $W_{1}$ satisfies the following upper bound:
$\left\|\frac{\partial y_{L}}{\partial W_{1}}\right\|_{2}\leq\prod_{\ell=1}^{L-1}\left(1+\frac{1}{\ell\sigma_{x_{\ell}}}A^{\prime}+\frac{1}{\ell^{2}\sigma_{x_{\ell}}^{2}}B^{\prime}\right),$ (18)
where $A^{\prime}$ and $B^{\prime}$ are dependent on the scaled neural network parameters defined in A.5.
From (18), we can easily get that if we do not want so many loss spikes, we need to let the $\left\|\frac{\partial y_{L}}{\partial W_{1}}\right\|_{2}$ do not explode. Which in our assumption means that the variance of the deep layer should not be too small. Based on the above result (15), the good variance growth rate is sub linearly growth. which is:
$\sigma^{2}_{x_{\ell}}\sim\ell,$ (19)
which is actually the LayerNorm Scaling convergence rate. Therefore, the LayerNorm Scaling method can provide a moderate scaling of the variance, both to make the deeper layers effective and to prevent the initial layers from exploding.
The proof of Theorem 4.3 is in Section A.5. Then we can easily generalize to a more general situation for layer $l$. By carefully controlling the propagation of gradients through the attention blocks, we can observe that for every layer $\ell$, the $\left\|\frac{\partial y_{L}}{\partial W_{\ell}}\right\|_{2}$ has the same upper bound as in Result 18. The proof is omitted here. As a result, LayerNorm Scaling improves stability for every layer (especially the first layer) and avoids sharp gradient amplification, which would otherwise result in an unstable or inefficient optimization process.
@@ 5.1 LLM Pre-training
To evaluate the effectiveness of LayerNorm Scaling, we follow the experimental setup of Li et al. (2024b), using the identical model configurations and training conditions to compare LNS with widely used normalization techniques, including Post-LN (Nguyen and Salazar, 2019), DeepNorm (Wang et al., 2024), and Pre-LN (Dai et al., 2019). In line with Lialin et al. (2023) and Zhao et al. (2024), we conduct experiments using LLaMA-based architectures with model sizes of 130M, 250M, 350M, and 1B parameters.
Table 1: Perplexity (↓) comparison of various layer normalization methods.
@@
The architecture incorporates RMSNorm (Shazeer, 2020) and SwiGLU activations (Zhang and Sennrich, 2019), which are applied consistently across all model sizes and normalization methods. For optimization, we use the Adam optimizer (Kingma, 2015) and adopt size-specific learning rates: $1\times 10^{-3}$ for models up to 350M parameters, and $5\times 10^{-4}$ for the 1B parameter model. All models share the same architecture, hyperparameters, and training schedule, with the only difference being the choice of normalization method. Unlike Mix-LN (Li et al., 2024b), which introduces an additional hyperparameter $\alpha$ manually set to 0.25, LayerNorm Scaling requires no extra hyperparameters, making it simpler to implement. Table 1 shows that LayerNorm Scaling consistently outperforms other normalization methods across different model sizes. While DeepNorm performs comparably to Pre-LN on smaller models, it struggles with larger architectures like LLaMA-1B, showing signs of instability and divergence in loss values. Similarly, Mix-LN outperforms Pre-LN in smaller models but faces convergence issues with LLaMA-350M, indicating its sensitivity to architecture design and hyperparameter tuning due to the introduction of Post-LN. Notably, Mix-LN was originally evaluated on LLaMA-1B with 50K steps (Li et al., 2024b), while our setting extends training to 100K steps, where Mix-LN fails to converge, highlighting its instability in large-scale settings caused by the usage of Post-LN.
In contrast, LayerNorm Scaling solves the Curse of Depth without compromising the training stability. LayerNorm Scaling achieves the lowest perplexity across all tested model sizes, showing stable performance improvements over existing methods. For instance, on LLaMA-130M and LLaMA-1B, LayerNorm Scaling reduces perplexity by 0.97 and 1.31, respectively, compared to Pre-LN. Notably, LayerNorm Scaling maintains stable training dynamics for LLaMA-1B, a model size where Mix-LN fails to converge. These findings demonstrate that LayerNorm Scaling provides a robust and computationally efficient normalization strategy, enhancing large-scale training of language models without additional implementation complexity.
@@ 5.2 Supervised Fine-tuning
To verify whether the gains in pre-training can be translated to the stage of post-training, we perform SFT with the models obtained from Section 5.1 on the Commonsense170K dataset Hu et al. (2023) across eight downstream tasks. We adopt the same fine-tuning configurations as used in Li et al. (2024b). The results, presented in Table 2, demonstrate that LayerNorm Scaling consistently surpasses other normalization techniques in all evaluated datasets. For the LLaMA-250M model, LayerNorm Scaling improves average performance by 1.80% and achieves a 3.56% gain on ARC-e compared to Mix-LN. Similar trends are observed with the LLaMA-1B model, where LayerNorm Scaling outperforms Pre-LN, Post-LN, Mix-LN, and DeepNorm on seven out of eight tasks, with an average gain of 1.86% over the best baseline. These results confirm that LayerNorm Scaling enhances generalization on diverse downstream tasks by improving the representation quality of deep layers.
Table 2: Fine-tuning performance ($\uparrow$) of LLaMA with various layer normalizations.
@@
5.3 Scaling Up Training
5.3.1 OLMo
Model Size Scaling. To further assess the scalability and robustness of LNS, we conduct additional experiments using the OLMo repository (Groeneveld et al., 2024), scaling training across model sizes of 60M, 150M, 300M, 1B, and 7B parameters. All models are trained on a fixed 20B-token budget to ensure comparability. These experiments are designed to evaluate whether the performance gains observed with LNS in smaller-scale settings extend to more challenging and state-of-the-art LLM training regimes. As shown in Figure 1, LNS consistently and substantially outperforms the standard Pre-LN baseline across all model sizes. Remarkably, for the 7B model, LNS reduces the final loss from 2.69 to 2.50. These results underscore the scalability of LNS and its effectiveness in large-scale pre-training scenarios.
Figure 5: Training loss of OLMo-7B with Pre-LN and LNS.
Loss Curve. Figure 5 shows the training loss curves of 7B models trained with Pre-LN and LNS. While LayerNorm Scaling exhibits slightly slower convergence at the early stages of training, it consistently outperforms Pre-LN as training progresses, ultimately achieving a substantial loss gap. We attribute this to the uncontrolled accumulation of output variance in Pre-LN, which amplifies with depth and training steps, ultimately impairing the effective learning of deeper layers. In contrast, LNS mitigates this issue by scaling down the output variance in proportion to depth, thereby enabling more stable and effective training across all layers during training.
Beating OLMo’s Scaled Initialization. OLMo adopts the scaled initialization proposed in Zhang et al. (2019) and used by Mehta et al. (2024), which scales input projections by $1/\sqrt{d_{\mathrm{model}}}$, and output projections by $1/\sqrt{2\cdot d_{\mathrm{model}}\cdot l}$ at every layer. This method is designed to enhance training stability and to scale down variance at initialization. To evaluate the effectiveness of LNS, we compare it against this state-of-the-art initialization by training OLMo-1B on 20B tokens. As shown in Table 3, LNS achieves consistently lower training loss, indicating that it may offer a more effective alternative for large-scale LLM training.
Table 3: Comparison with OLMo’s Scaled Initialization.
Method Model # Tokens Training Loss Perplexity
OLMo’s Scaled Initialization OLMo-1B 20B 2.96 19.30
LayerNorm Scaling OLMo-1B 20B 2.85 17.28
5.3.2 Qwen2.5
We further evaluate the generalizability of LNS by applying it to a state-of-the-art architecture, Qwen2.5-0.5B (Yang et al., 2024). We train the model for 6B tokens and compare LNS against the standard Pre-LN setup. Consistent with previous findings, Table 4 illustrates that LNS yields a notable reduction in perplexity—from 20.62 to 19.57—highlighting its effectiveness even on strong, modern architectures.
Table 4: Perplexity (PPL $\downarrow$) comparison under scaled-up pre-training. For LLaMA-1B and 7B, training is scheduled for 100B tokens but is terminated early to report results. Qwen-2.5 is trained with a fixed budget of 6B tokens.
Model # Params # Tokens Pre-LN (PPL) LNS (PPL)
Qwen2.5-0.5B 0.5B 6B 20.62 19.57
The consistent benefits observed across increased model scales, larger training datasets, and diverse architectures suggest that LNS is a promising technique for enhancing the training of contemporary large language models, ensuring that deeper layers contribute more effectively to learning.
5.4 LNS Effectively Scales Down Output Variance
As LNS is proposed to reduce output variance, we empirically validate this claim during the pre-training of LLMs. We compare the layerwise output variance of three configurations: (1) the standard Pre-LN (Ba, 2016), (2) Pre-LN with Scaled Initialization (Shoeybi et al., 2020; Radford et al., 2019), which scales the initialization of the feedforward layers’ weights $W_{0}$ and $W_{2}$ by $\frac{1}{\sqrt{2L}}$, where $L$ is the total number of Transformer layers, and (3) Pre-LN with LNS. The average output variance across layers is shown in Figure 4. For both vanilla Pre-LN and Scaled Initialization, the output variance in shallow layers (blue) remains relatively stable throughout training, while variance in deeper layers (red) grows substantially after 2K iterations, reaching up to 175 in the final layer. Since Scaled Initialization only operates at initialization, it is insufficient to constrain output variance during training. In contrast, LNS consistently suppresses the growth of output variance in deeper layers, capping it at approximately 25.
5.5 LNS Enhances the Effectiveness of Deep Layers
Figure 6: Left: Performance drop of layer pruning on LLaMA-130M. Right: The angular distance between representations of subsequent layers is shown. LayerNorm Scaling enables deep layers to make a meaningful contribution to the model.
Furthermore, to assess whether LNS enhances the effectiveness of deeper layers by promoting more diverse feature representations, we analyze the layerwise performance drop and the angular distance of LNS, as shown in Figure 6. Compared to Pre-LN, the performance degradation in LayerNorm Scaling is more uniformly distributed across layers, indicating a more balanced contribution from each layer. Notably, pruning the deeper layers of LNS results in a more significant accuracy drop, suggesting these layers play a more critical role in task performance. Additionally, features learned under LNS exhibit greater distinction: most layers show a substantial angular distance, exceeding 0.6, from their adjacent layers, indicating more diverse representations. In sharp contrast, the layerwise angular distance in Pre-LN remains significantly lower and progressively decreases with depth, suggesting reduced feature diversity.
5.6 LayerNorm Scaling in Vision Transformer
To evaluate whether LNS also benefits architectures beyond language models, we conduct experiments on ViT-S on ImageNet-1K. Since ViT-S includes LayerScale (Touvron et al., 2021) by default—which may interfere with the effect of LNS—we remove LayerScale from all evaluated variants to ensure a fair comparison. We then test different insertion positions of LNS. The top-1 accuracy results are summarized in Table 5. Whereas LNS in language models is typically most effective directly after normalization, in Vision Transformers, the best position is after the attention and MLP blocks. We next examine whether this performance gain correlates with better control of layer-wise variance.
Table 5: Top-1 accuracy (%) of ViT-S model with and without LNS.
Model Variant LNS Position Top-1 Accuracy
ViT (w/o LayerScale) 67.91
ViT (w/o LayerScale) after LayerNorm 66.43
ViT (w/o LayerScale) after Attn/MLP 68.75
Figure 7 plots the average output variance of each transformer block during training. Without LayerScale, variance in deeper layers grows rapidly—exceeding $\sim\!3{,}000$ by 30K update steps. Applying LNS after Attn/MLP controls this growth to below $\sim\!150$, confirming that LNS stabilizes the forward signal even in vision transformers.
Figure 7: Layer-wise output variance of ViT-S without LayerScale (left) and with LNS after Attn/MLP (right). LNS significantly reduces the variance growth compared to the baseline.
These preliminary findings indicate that the variance-control mechanism underlying LNS generalizes to vision transformers when the scaling is applied after Attn/MLP. We leave a more detailed theoretical understanding of this behavior to future work and community discussion.
6 Ablation Study
Comparing Against Other Scaling Methods. We first compare LNS with previous scaling approaches, including (1) Scaled Initialization (Shoeybi et al., 2020; Radford et al., 2019), which scales the initialization of $W_{0}$ and $W_{2}$ by the overall depth $1/\sqrt{2L}$; (2) Depth-Scaled Initialization (Zhang et al., 2019) scales the initialization of weight matrices by the current depth $1/\sqrt{2l}$; (3) SkipInit (De and Smith, 2020) introduces a learnable parameter after FFN/Att layers, initialized as $1/\sqrt{L}$; (4) LayerScale (Touvron et al., 2021) applies per-channel weighting using a diagonal matrix, diag$(\lambda_{1},\ldots,\lambda_{d})$, where each weight $\lambda_{i}$ is initialized to a small value (e.g., $\lambda_{i}=\epsilon$). Table 6 presents the results of LLaMA-130M and LLaMA-250M.
First, we observe that methods involving learnable parameters, such as LayerScale and SkipInit, consistently degrade performance in LLMs. Among initialization-based techniques, a larger scaling factor proves beneficial: Scaled Initialization yields lower perplexity compared to Depth-Scaled Initialization. Notably, LNS achieves the best overall performance, underscoring the advantage of applying scaling dynamically during training. Interestingly, combining LNS with Scaled Initialization results in worse performance than using LNS alone, highlighting the importance of removing conflicting initialization strategies prior to adopting LNS.
@@ 6 Ablation Study
Table 6: Comparing LNS against other scaling methods. Perplexity (↓) is reported.
LLaMA-130M LLaMA-250M
@@
+ LayerScale (Touvron et al., 2021) 27.93 23.45
+ SkipInit (De and Smith, 2020) 27.41 22.29
+ Depth-Scaled Initialization (Zhang et al., 2019) 26.95 21.50
+ Scaled Initialization (Shoeybi et al., 2020) 26.04 20.98
@@
+ LayerNorm Scaling + Scaled Initialization 25.80 20.79
Comparison with Other Layer Normalization. In addition, we conducted comparisons using LLaMA-130M to evaluate LayerNorm Scaling against recently proposed normalization methods, including Admin (Liu et al., 2020), Sandwich-LN (Ding et al., 2021), Group-LN (Wu and He, 2018; Ma et al., 2024), and Mix-LN (Li et al., 2024b). Table 7 shows that Admin and Group-LN degrade performance. Sandwich-LN slightly outperforms Pre-LN. Both Mix-LN and LayerNorm Scaling improve over Pre-LN by good margins. However, Mix-LN fails to reduce perplexity under 26, falling short of LayerNorm Scaling and suffers from instability in large-scale scenarios as shown in Table 1.
Table 7: Comparison against other normalization methods on LLaMA-130M. All methods use the identical configurations. Perplexity (↓) is reported.
Pre-LN Admin Group-LN Sandwich-LN Mix-LN LayerNorm Scaling
26.73 27.91 28.01 26.51 26.07 25.76
Effect of Positions of LNS. The results in Table 8 show that inserting the scaling factor at different points can have a considerable influence on the model’s performance. Placing it after the residual connection (“After Residual”) leads to a perplexity of 1358.11, which indicates training divergence. In contrast, LNS incorporates the scaling factor after LN achieving the best perplexity of 25.76, surpassing both the baseline Pre-LN setting (26.73) and other placements. This suggests that modifying the LayerNorm to include the scaling factor can enhance training stability and final performance for this model configuration.
Table 8: Effects of Insertion Position of LayerNorm Scaling on LLaMA-130M
Pre-LN Before LN After Attn/FFN After Residual LNS Only After Attn LNS Only After FFN Ours (After LN)
26.73 26.97 26.53 1358.11 26.89 26.43 25.76